### v1

# Homotopy Methods for Convex Optimization

Abstract
Convex optimization encompasses a wide range of optimization problems, containing many efficiently solvable subclasses. Interior point methods are currently the state-of-the-art approach for solving such problems, particularly effective for classes like semidefinite programming, quadratic programming, and geometric programming. However, their success hinges on the construction of self-concordant barrier functions for the feasible sets. In this work, we introduce an alternative method for tackling convex optimization problems, employing a homotopy. With this technique, the feasible set of a trivial optimization problem is continuously transformed into the target one, while tracking the solutions. We conduct an analysis of this approach, focusing on its application to semidefinite programs, hyperbolic programs, and convex optimization problems with a single convexity constraint. Moreover, we demonstrate that our approach numerically outperforms state-of-the-art methods in several interesting cases.
1 Introduction
Optimization problems with convex constraints and a convex objective function (aka convex optimization problems) are a class of optimization problems that find applications in quantum information theory, portfolio optimization, data fitting, and many more [5]. The state-of-the-art methods to solve these problems are interior-point methods, that go back to Nesterov and Nemirovski [21]. While these are successfully applied to many sub-classes of convex optimization problems, including semidefinite programming, quadratically constrained quadratic programs, or geometric programming, they don’t solve arbitrary convex optimization problems efficiently. The success of these methods heavily rely on the construction of a self-concordant barrier function.
In this paper, we present a way to solve certain convex optimization problems using a homotopy method. We apply this method to convex optimization problems with a single convexity constraint, as well as semidefinite and hyperbolic programs. The presented method, as all homotopy methods, is based on the following general idea:
Homotopically change a problem with obvious solutions into the target problem and follow the path of solutions along the homotopy.
More precisely, we start from a convex optimization problem with an obvious solution (i.e. analytically solvable) and homotopically change it to the target problem (see Fig. 1). In contrast to some other homotopy methods, we propose to change the feasible set of the optimization problem, instead of changing the objective function. We then show that the path of optimal solutions is the unique solution of a system of differential equations, and we numerically solve this system to obtain a solution of the optimization problem.
$\mathbf{b}$$\mathbf{a}_{0}$$\mathbf{b}$$\mathbf{a}_{t}$$\mathbf{b}$$\mathbf{a}_{1}$
Figure 1: Visualization of the method presented in this paper. To solve the convex optimization problem on the right (i.e. maximizing $\mathbf{x}\mapsto\mathbf{b}^{t}\mathbf{x}$ over a convex set), one instead solves a trivial convex optimization problem (left) and tracks the path of optimal solutions via solving a differential equation.
Main contributions.
We introduce the notion of a smooth and regular description of families of convex sets, and show that the path of optimal solutions is the unique solution of a certain differential equation (see Theorem 6 for the main result of this paper). We then construct explicit homotopies for two optimization classes and show that these homotopies are smooth and regular. These examples include:
  1. (i)
    Semidefinite and Hyperbolic programming.
  2. (ii)
    Optimization problems with a linear objective and one convexity constraint.
We further benchmark the homotopy method in the following four examples:
  1. (a)
    Optimizing on the hyperbolicity cone of elementary symmetric polynomials.
  2. (b)
    Optimizing a linear objective function over a $k$-ellipsoid.
  3. (c)
    Optimizing a linear objective function over the $p$-norm ball.
  4. (d)
    Geometric programming with a single constraint.
In particular, we show that the homotopy method outperforms the lifting techniques to semidefinite programs in (a), (b), and (c).
Homotopic approaches in other fields.
The idea of homotopy methods has been successfully applied in various different fields, including algebraic geometry, non-linear partial differential equations, and quantum computing:
  • $\triangleright$
    In numerical algebraic geometry (originally developed by Sommese and Wampler [31], see [11, 32] for an overview) the main objective is to solve a system of polynomial equations
    $p_{1}(\mathbf{z})=0,\ldots,p_{k}(\mathbf{z})=0$
    in multiple variables $\mathbf{z}\in\mathbb{C}^{n}$ with finitely many solutions. First, a system of polynomial equations
    $g_{1}(\mathbf{z})=0,\ldots,g_{k}(\mathbf{z})=0$
    with trivial solutions is generated. This could be for example the unit polynomials $g_{i}(\mathbf{z})=z_{i}^{d}$. Afterwards, these solutions are tracked via the homotopy
    $(1-t)g_{i}(\mathbf{z})+tp_{i}(\mathbf{z})=0\text{ for every }i\in\{1,\ldots,k\}.$
    More specifically, the homotopy method consists of letting evolve the parameter $t$ from 00 to $1$ and following the path of solutions along $t$ via solving a differential equation.
  • $\triangleright$
    The homotopy analysis method [17, 19, 18] and the homotopy perturbation method [13] are approaches to solve non-linear partial differential equation via infinitely many linear ones. By setting up a homotopy between a linear partial differential operator $\mathcal{L}$ to a non-linear one $\mathcal{N}$, one obtains a homotopy of solutions $(u_{t})_{t\in[0,1]}$, where $u_{0}$ is the solution of the initial linear equation $\mathcal{L}$ and $u_{1}$ is the solution of the target non-linear equation $\mathcal{N}$. Via a Taylor expansion
    $u_{t}(\mathbf{x})=u_{0}(\mathbf{x})+\sum_{k=1}^{\infty}v_{k}(\mathbf{x})t^{k}$
    one obtains linear differential equations for every $v_{k}$.
  • $\triangleright$
    In adiabatic quantum computing and quantum annealing [9], the main objective is to compute the ground state of a quantum-Hamiltonian $H_{B}$. This is in general an $\mathsf{NP}$-hard optimization problem [3]. Let
    $H_{t}\coloneqq(1-t)\cdot H_{A}+t\cdot H_{B}$
    describe a physical system which is described at time $t=0$ by a Hamiltonian $H_{A}$ with a trivial ground state and at time $t=1$ by the target Hamiltonian $H_{B}$. If the evolution from $H_{A}$ to $H_{B}$ happens slowly enough, then it is guaranteed that the state of the system remains in its ground state for every $t,$ which implies that one can obtain the ground state of $H_{B}$ when knowing the ground state of $H_{A}$. Since ground-states of quantum Hamiltonians can embed $\mathsf{NP}$-complete optimization problems, adiabatic quantum computing can be employed as a homotopy method solving $\mathsf{NP}$-complete optimization problems [20].
Related work.
So far, homotopy methods have also been applied in semidefinite programming in different contexts, which are however different from what we propose in this paper:
  • $\triangleright$
    To apply the interior point method to semidefinite programs, they have to be strictly feasible. By introducing a perturbation on the feasible set, this condition can be relaxed [12, 15].
  • $\triangleright$
    Convex optimization problems with algebraic constraints can be solved via a system of polynomial equations arising from the Karush–Kuhn–Tucker (KKT) conditions (see for example [5]). This system can be solved via numerical algebraic geometry (see for example [6] for an implementation and [28] for a related problem).
Overview.
This paper is structured as follows: In Section 2 we introduce the notion of a homotopy between optimization problems, and then show under which conditions the induced differential equation of optimal solutions has a unique solution. In Section 3 we apply our homotopy method to two optimization classes: hyperbolic programs (which include semidefinite programs) and convex optimization problems with a single constraint function. In Section 4 we provide some numerical examples that compare our method to existing methods. In Section 5 we present open questions and further directions. Finally, we present in Appendix A another canonical homotopy between semidefinite programs which does not behave well regarding our method, and in Appendix B we show that the method also generalizes to quasi-concave objective functions $f$ and when introducing a homotopy on $f$. Appendix C contains the proof of boundedness of rigidly convex sets defined by elementary symmetric polynomials, used in Section 4.
2 Homotopies Between Convex Optimization Problems
In the following, we introduce the optimization problem we want to solve. Moreover, we introduce the notion of a homotopy between optimization problems and the ordinary differential equation for the optimal points. This section is structured as follows:
  • $\triangleright$
    In Section 2.1 we introduce the notion of a homotopy between optimization problems.
  • $\triangleright$
    In Section 2.2 we derive the initial value problem for stationary points of the optimization problem.
  • $\triangleright$
    In Section 2.3 we prove that the solution of the initial value problem gives rise to the optimal value of the optimization problem. This gives rise to the main theoretical statement of this paper in Theorem 6.
2.1 Homotopies of convex sets and problem statement
Let $f\colon\mathbb{R}^{n}\to\mathbb{R}$ be a linear function, and let $C\subseteq\mathbb{R}^{n}$ be a compact convex set. Our objective is to find an optimal point (and thus the optimal value) of the optimization problem
$\max_{\mathbf{x}\in C}\ f(\mathbf{x}).$
We do this by continuously deforming the feasible set from an easy one into $C$, and keeping track of an optimal solution of the corresponding optimization problems along the way. Note that this construction can be generalized to concave functions $f$, by introducing a homotopy of objective functions. In Appendix B we elaborate on this more general case.
To be able to define and work with the homotopies, we will assume that the feasible sets come with a description as follows:
Definition 1 (Smooth description).
A smooth description for the compact convex set $C\subseteq\mathbb{R}^{n}$ is a continuously differentiable function $p\colon\mathbb{R}^{n}\to\mathbb{R}$ with
$p(\mathbf{x})=0\ \mbox{ and }\ \nabla p(\mathbf{x})\neq 0\ \mbox{ for all }\mathbf{x}\in\partial C.$
Capturing the boundary of $C$ is enough, since the maximum of $f$ is always attained on $\partial C$. However, note that we do not require $p(\mathbf{x})=0$ to imply $\mathbf{x}\in\partial C$ (see for example real zero polynomials in Section 3.1). In particular, the set $C$ is in general not uniquely determined by a smooth description, so we still have to keep track of it. Also note that the existence of a smooth description imposes conditions on the set $C$. For example, it implies that $C$ is smooth, in the sense that each boundary point admits a unique supporting hyperplane.
We say that a twice continuously differentiable function $p\colon\mathbb{R}^{n}\to\mathbb{R}$ is concave at a point $\mathbf{x}\in\mathbb{R}^{n}$, if its Hessian matrix $\nabla^{2}p(\mathbf{x})$ is negative semidefinite, i.e. all its eigenvalues are non-positive. It is strictly concave if $\nabla^{2}p(\mathbf{x})$ is negative definite, i.e. all eigenvalues are strictly negative. We say that $p$ is strictly quasi-concave at $\mathbf{x}\in\mathbb{R}^{n}$, if the quadratic form $\mathbf{v}\mapsto\mathbf{v}^{t}\nabla^{2}p(\mathbf{x})\mathbf{v}$ is negative definite on the orthogonal complement of $\nabla p(\mathbf{x})$. In other words,
$\mathbf{v}^{t}\nabla p(\mathbf{x})=0\quad\Longrightarrow\quad\mathbf{v}^{t}\nabla^{2}p(\mathbf{x})\mathbf{v}<0.$ (1)
for every $\mathbf{v}\in\mathbb{R}^{n}\setminus\{0\}$. [^1] This definition is stronger than the usual definition of strict (quasi-)concavity (see for example Section 3.4.2 and Section 3.4.3 of [5]). However, all (quasi-)concave functions in this paper satisfy this stronger type of (quasi-)concavity.
If a smooth description of $C$ is strictly quasi-concave at a point $\mathbf{x}\in\partial C$, the point is an exposed point, i.e. the unique supporting hyperplane at $\mathbf{x}$ intersects $C$ only in $\mathbf{x}$. In particular, if this holds for all boundary points, then maximizing a linear function over $C$ has a unique solution.
Now let $\left(C_{t}\right)_{t\in[0,1]}$ be a family of compact convex subsets of $\mathbb{R}^{n}.$ For our algorithm we need a suitable homotopy of smooth descriptions.
Definition 2.
(i) A homotopy of smooth descriptions for the family $\left(C_{t}\right)_{t\in[0,1]}$ is a continuously differentiable function
$p\colon[0,1]\times\mathbb{R}^{n};\ (t,\mathbf{x})\mapsto p_{t}(\mathbf{x})$
such that $p_{t}$ provides a smooth description of $C_{t}$ for all $t\in[0,1]$.
(ii) The homotopy $p$ of smooth descriptions is called regular, if for all $t^{*}\in[0,1],$ $\varepsilon>0,$ and all continuously differentiable functions $\mathbf{a}\colon[t^{*}-\varepsilon,t^{*}+\varepsilon]\to\mathbb{R}^{n}$ with $\mathbf{a}_{t^{*}}\in\partial C_{t^{*}}$ and $p_{t}(\mathbf{a}_{t})=0$ for all $t\in[t^{*}-\varepsilon,t^{*}+\varepsilon],$ we have $\mathbf{a}_{t}\in\partial C_{t}$ for every $t\in[t^{*}-\varepsilon,t^{*}+\varepsilon]$. Note that we write $\mathbf{a}_{t}$ for $\mathbf{a}(t)$.
Remark 3.
In words, regularity means that we cannot leave the boundary of the sets over time, without the functions $p_{t}$ noticing. Not every homotopy of smooth descriptions is regular. As a counterexample one can construct a homotopy of smooth descriptions for a family that switches in between two disjoint balls.
But under a certain continuity condition on the family $\left(C_{t}\right)_{t\in[0,1]}$, every homotopy of smooth descriptions in indeed regular. Assume, for example, that
$B:=\bigcup_{t\in[0,1]}\partial C_{t}\times\{t\}\subseteq\mathbb{R}^{n+1}$
is an $n$-dimensional manifold (with boundary $\partial C_{0}\cup\partial C_{1}$). Then at each point $(\mathbf{x},t)\in B$ the gradient of $p$ is nonzero, already in space direction, since $p_{t}$ is a smooth description of $C_{t}$. Thus the implicit function theorem implies that the zero set of $p$ is the graph of a function in $t$ and $n-1$ of the space variables, locally around $(\mathbf{x},t)$. But since $B$ is an $n$-dimensional manifold contained in this zero set/graph, it coincides with it locally.
Now for a continuous function $\mathbf{a}\colon[t^{*}-\varepsilon,t^{*}+\varepsilon]\to\mathbb{R}^{n}$ with $p_{t}(\mathbf{a}_{t})=0$ for all $t$, the set
$I=\{t\in[t^{*}-\varepsilon,t^{*}+\varepsilon]\mid\mathbf{a}_{t}\in\partial C_{t}\}$
is easily checked to be open and closed. From $\mathbf{a}_{t^{*}}\in\partial C_{t^{*}}$ we obtain $t^{*}\in I$ and thus $I=[t^{*}-\varepsilon,t^{*}+\varepsilon].$
In the context of the homotopy method, we now consider for every $t\in[0,1]$ the optimization problem
$\max_{\mathbf{x}\in C_{t}}\ f(\mathbf{x})$ (2)
and we denote an optimal point by $\mathbf{a}_{t}\in\mathbb{R}^{n}$. To solve the optimization problems (2), we will make use of the fact that the condition $p_{t}(\mathbf{x})=0$ is necessary to have that $\mathbf{x}\in\partial C_{t}$. For this reason, we instead consider the following optimization problem
$\begin{split}\max_{\mathbf{x}\in\mathbb{R}^{n}}\ &f(\mathbf{x})\\ \text{ subject to }&p_{t}(\mathbf{x})=0\end{split}$ (3)
and use that fact that any optimal point of (2) is a stationary point of (3). We will then obtain a differential equation whose solution is a function $\mathbf{a}\colon[0,1]\to\mathbb{R}^{n}$ of stationary points of (3). We prove in Theorem 6 that this path of stationary points are the optimal points of (2).
2.2 The Differential Equation for Stationary Points
We now derive a condition for stationary points of the optimization problem (3) for every $t\in[0,1]$ via Lagrange multipliers. Afterwards, we translate these conditions into an initial value problem (see Eq. 6).
Using Lagrange multipliers, one obtains that $\mathbf{a}_{t}$ is a stationary point222i.e. $f(\mathbf{a}_{t})$ is either a local minimum, local maximum or a saddle point when considered on the set $\{\mathbf{x}\in\mathbb{R}^{n}\mid p_{t}(\mathbf{x})=0\}$. of the problem (3) at time $t$ (assuming $\nabla p_{t}(\mathbf{a}_{t})\neq 0$) if and only if
  1. ($\text{C}_{1}$)
    $\nabla f(\mathbf{a}_{t})=\lambda_{t}\nabla p_{t}(\mathbf{a}_{t})$
  2. ($\text{C}_{2}$)
    $p_{t}(\mathbf{a}_{t})=0$
where $\lambda_{t}\in\mathbb{R}$ is the Lagrange multiplier for the constraint $p_{t}$.
In the following, we translate these two conditions into an initial value problem, whose solution is a path $\mathbf{a}\colon[0,1]\to\mathbb{R}^{n};t\mapsto\mathbf{a}_{t}$ where $\mathbf{a}_{t}$ is a stationary point of the constrained optimization problem at time $t$.
We first formulate an equivalent condition for ($\text{C}_{1}$) using the following Lemma.
Lemma 4.
Let $\mathbf{v},\mathbf{w}\in\mathbb{R}^{n}\setminus\{0\}$. If $v_{k}\neq 0$, we have that $\mathbf{v}=\lambda\mathbf{w}$ for some $\lambda\in\mathbb{R}$ if and only if
$v_{k}w_{i}-v_{i}w_{k}=0$ (4)
for every $i\in\{1,\ldots,n\}$.
Proof.
The only if direction is immediate. So let $\mathbf{v},\mathbf{w}\in\mathbb{R}^{n}\setminus\{0\}$ such that (4) holds. Note that also $w_{k}\neq 0$ holds, since otherwise we would get $\mathbf{w}=0.$ Now we define $\lambda$ via $v_{k}=\lambda w_{k}$. We now show that $v_{i}=\lambda w_{i}$ for every $i$. We have two cases:
  1. (i)
    $w_{i}\neq 0$. Then $v_{i},w_{k}\neq 0$ and we have that
    $\frac{v_{i}}{w_{i}}=\frac{v_{k}}{w_{k}}=\lambda$
  2. (ii)
    $w_{i}=0$. Then $v_{i}w_{k}=0$. Since $w_{k}\neq 0$, we have that $v_{i}=0$ which implies that $v_{i}=\lambda w_{i}$.∎
By Lemma 4, ($\text{C}_{1}$) is equivalent to
$\mathcal{Q}_{k,i}(\mathbf{a}_{t},t)\coloneqq\partial_{k}f(\mathbf{a}_{t})\partial_{i}p_{t}(\mathbf{a}_{t})-\partial_{i}f(\mathbf{a}_{t})\partial_{k}p_{t}(\mathbf{a}_{t})=0$ (5)
[…]
Differentiating (5) and ($\text{C}_{2}$), we obtain an initial value problem for stationary points $\mathbf{a}\colon[0,1]\to\mathbb{R}^{n}$:
[…]
00 0 $=\big{[}\nabla\mathcal{Q}_{k,i}(\mathbf{a}_{t},t)\big{]}^{t}\cdot\frac{\mathrm{d}\mathbf{a}_{t}}{\mathrm{d}t}+\partial_{t}\mathcal{Q}_{k,i}(\mathbf{a}_{t},t)\quad\text{ for }i\in\{1,\ldots,k-1,k+1,\ldots,n\}$
00 0 $=\big{[}\nabla p_{t}(\mathbf{a}_{t})\big{]}^{t}\cdot\frac{\mathrm{d}\mathbf{a}_{t}}{\mathrm{d}t}+\partial_{t}p_{t}(\mathbf{a})$
where $k$ is chosen such that $\partial_{k}f\neq 0$. In summary, for every solution $\mathbf{a}\colon t\mapsto\mathbf{a}_{t}$ we have that $\mathbf{a}_{t}$ is a stationary point of the constraint optimization problem of $p_{t}$ and $f$. Equivalently, the differential equation reads as
$K(\mathbf{a}_{t},t)\cdot\frac{\mathrm{d}\mathbf{a}_{t}}{\mathrm{d}t}+m(\mathbf{a}_{t},t)=0$ (6)
where
$K(\mathbf{x},t)=\begin{bmatrix}\nabla p_{t}(\mathbf{x})^{t}\\ \nabla\mathcal{Q}_{k,1}(\mathbf{x},t)^{t}\\ \vdots\\ \nabla\mathcal{Q}_{k,k-1}(\mathbf{x},t)^{t}\\ \nabla\mathcal{Q}_{k,k+1}(\mathbf{x},t)^{t}\\ \vdots\\ \nabla\mathcal{Q}_{k,n}(\mathbf{x},t)^{t}\end{bmatrix}\quad\text{ and }\quad m(\mathbf{x},t)=\begin{bmatrix}\partial_{t}p_{t}(\mathbf{x})\\ \partial_{t}\mathcal{Q}_{k,1}(\mathbf{x},t)\\ \vdots\\ \partial_{t}\mathcal{Q}_{k,k-1}(\mathbf{x},t)\\ \partial_{t}\mathcal{Q}_{k,k+1}(\mathbf{x},t)\\ \vdots\\ \partial_{t}\mathcal{Q}_{k,n}(\mathbf{x},t)\end{bmatrix}.$ (7)
2.3 Existence of a Unique Path of Optimal Solutions
We now prove that there is a unique solution for the differential equation (6), if $\mathbf{x}\mapsto p_{t}(\mathbf{x})$ is a strictly quasi-concave function for every $t\in[0,1)$. [^4] This means that the target convex set $C_{1}$ does not need to be strictly quasi-convex for the success of the method. The proof is split into two parts:
  1. (i)
    In Lemma 5 we show that $K(\mathbf{a}_{t},t)$ is invertible, if $\mathbf{a}_{t}$ is a stationary point at time $t$.
  2. (ii)
    In Theorem 6, we prove the solution is unique if $K(\mathbf{a}_{t},t)$ is invertible for every $t$.
We now show that $K(\mathbf{a}_{t},t)$ is invertible under certain assumptions on $p_{t}$.
Lemma 5.
Let $\mathbf{a}_{t}$ be a stationary point of the constrained problem (3) at time $t$. If $p_{t}$ is strictly quasi-concave at $\mathbf{a}_{t}$, then $K(\mathbf{a}_{t},t)$ is invertible.
Proof.
We show that for every $0\neq\mathbf{v}\in\mathbb{R}^{n}$, we have that
$\mathbf{v}^{t}\nabla p_{t}(\mathbf{a}_{t})\neq 0\quad\text{ or }\quad\mathbf{v}^{t}\nabla\mathcal{Q}_{k,i}(\mathbf{a}_{t},t)\neq 0\text{ for some }i\in\{1,\ldots,k-1,k+1,\ldots,n\}.$
So assume that $\mathbf{v}^{t}\nabla p_{t}(\mathbf{a}_{t})=0$. Since $\mathbf{a}_{t}$ is a stationary point we have $\mathbf{v}^{t}\nabla f(\mathbf{a}_{t})=0$, and by strict quasi-concavity of $p_{t}$ that $\mathbf{v}^{t}\nabla^{2}p_{t}(\mathbf{a}_{t})\mathbf{v}\neq 0.$ For every $i\in\{1,\ldots,k-1,k+1,\ldots,n\}$, we can rewrite
$\nabla\mathcal{Q}_{k,i}(\mathbf{a}_{t},t)^{t}\mathbf{v}=\partial_{k}f(\mathbf{a}_{t})\big{(}\nabla^{2}p_{t}(\mathbf{a}_{t})\mathbf{v}\big{)}_{i}-\partial_{i}f(\mathbf{a}_{t})\big{(}\nabla^{2}p_{t}(\mathbf{a}_{t})\mathbf{v}\big{)}_{k}$
By Lemma 4, we have that $\nabla\mathcal{Q}_{k,i}(\mathbf{a}_{t},t)^{t}\mathbf{v}=0$ for all $i\in\{1,\ldots,k-1,k+1,\ldots,n\}$ if and only if $\nabla^{2}p_{t}(\mathbf{a}_{t})\mathbf{v}$ is parallel to $\nabla f(\mathbf{a}_{t})$. But since we have
$\mathbf{v}^{t}\nabla^{2}p_{t}(\mathbf{a}_{t})\mathbf{v}\neq 0\quad\text{ and }\quad\mathbf{v}^{t}\nabla f(\mathbf{a}_{t})=0$
we get a contradiction. ∎
We are now ready to show that the initial value problem (6) has a unique solution, which gives us the solutions of (2). This is the main theoretical result of our paper.
Theorem 6.
Let $\left(C_{t}\right)_{t\in[0,1]}$ be a family of compact convex sets, all with nonempty interior in $\mathbb{R}^{n}$. Let $p$ be a homotopy of smooth descriptions satisfying the following properties:
  • $\triangleright$
    $p$ is regular
  • $\triangleright$
    $p$ is three times continuously differentiable
  • $\triangleright$
    $p_{t}$ is strictly quasi-concave on $\partial C_{t}$ for all $t\in[0,1)$.
Then, if the initial value $\mathbf{a}_{0}$ is an optimal point of the optimization problem (2) for $t=0$, the initial value problem (6) has a unique solution, which consists of the optimal points of the problems (2).
Proof.
For $t^{*}\in[0,1)$ let $\mathbf{a}_{t^{*}}\in\partial C_{t^{*}}$ be the optimal point of the optimization problem (2), where uniqueness follows from strict quasi-concavity. By Lemma 5 we know that $K(\mathbf{a}_{t^{*}},t^{*})$ is invertible. Since invertibility is an open condition, this is also the case for every element contained in $R\coloneqq B_{\varepsilon}(\mathbf{a}_{t^{*}})\times[t^{*}-\varepsilon,t^{*}+\varepsilon]$ for a suitable choice of $\varepsilon>0$. Hence, we can locally define the differential equation
$\frac{\mathrm{d}\mathbf{x}}{\mathrm{d}t}=\mathcal{F}(\mathbf{x},t)\coloneqq-K(\mathbf{x},t)^{-1}\cdot m(\mathbf{x},t).$
Since the entries of $m(\mathbf{x},t)$ and $K(\mathbf{x},t)^{-1}$ are still continuously differentiable, we have that $\mathcal{F}_{|_{R}}$ is Lipschitz-continuous on $R$. Hence there exists a unique solution $\mathbf{a}\colon[t^{*}-\varepsilon,t^{*}+\varepsilon]\to B_{\varepsilon}(\mathbf{a}_{t^{*}})$ with $\mathbf{a}(t^{*})=\mathbf{a}_{t^{*}}$, by the Picard–Lindelöf Theorem.
From regularity of $p$ we obtain $\mathbf{a}_{t}\in\partial C_{t}$ for all $t\in[t^{*}-\varepsilon,t^{*}+\varepsilon]$. This implies that all $\mathbf{a}_{t}$ are the optimal points of (2). In fact, since all $C_{t}$ have nonempty interior, and $\mathbf{a}$ is continuous, a jump from the optimal point to the least optimal point, which is the other critical point on the boundary, is not possible. Hence, the local solutions must glue to a global solution.
This gives rise to a unique solution $\mathbf{a}\colon[0,1)\to\mathbb{R}^{n}$. By the boundedness and continuity of $\mathbf{a}$, there is a unique limiting point
$\mathbf{a}_{1}\coloneqq\lim_{t\uparrow 1}\mathbf{a}_{t}$
which is the optimal solution of the target problem. ∎
3 Examples of Homotopies
In the following, we present two classes of examples of homotopies that fulfill the assumptions from Theorem 6 above. In Section 3.1 we apply our method to the class of hyperbolic and semidefinite programs. In Section 3.2 we apply our method to convex optimization problems, where the constraint is generated by a single concave function.
While the homotopy in Section 3.2 has the desired properties for basic reasons, the homotopy for hyperbolic and real zero polynomials is based on the smoothing theory of hyperbolic polynomials which goes back to Nuij [26] and Renegar [30]. For self-consistency reasons, we will state the most important results in Section 3.1.
3.1 Semidefinite and Hyperbolic Programming
We now consider the problem of maximizing a linear function on a spectrahedron with non-empty interior [^5] We assume again without loss of generality that $\mathbf{0}$ is an interior point of the set of feasible points., i.e.
$\max_{\mathbf{x}\in\mathbb{R}^{n}}$ $f(\mathbf{x})$
subject to $A(\mathbf{x})\succcurlyeq 0$
where
$A(\mathbf{x})=I_{s}+x_{1}A_{1}+\cdots+x_{n}A_{n}$
with $A_{1},\ldots,A_{n}\in\operatorname{Sym}_{s}(\mathbb{R})$ being real symmetric matrices of size $s\times s$, $I_{s}$ the identity matrix of size $s$, and $X\succcurlyeq 0$ means that $X$ is positive semidefinite. This convex optimization problem is known as a semidefinite program. The feasible set
$\mathcal{S}(A)\coloneqq\{\mathbf{x}\in\mathbb{R}^{n}\mid A(\mathbf{x})\succcurlyeq 0\}$
is called aspectrahedron. [^6] Note that in general a spectrahedron is defined using an arbitrary symmetric matrix $A_{0}$ instead of $I_{s}$. However, every spectrahedron with non-empty interior can be represented (up to a translation) in this form (see for example Lemma 2.12 and Corollary 2.16 in [23]).
Generalizing spectrahedra to rigidly convex sets (see Definition 7) leads to the notion of hyperbolic programming. Rigidly convex sets arise from so-called real zero polynomials and share many properties with spectrahedra. The generalized Lax conjecture actually postulates that every rigidly convex set is a spectrahedron. While this statement is true for $n=2$, it is open in higher dimensions [16].
Note that our homotopy approach is useful beyond the established methods for solving semidefinite programs. First, as just mentioned, it is not clear whether each rigidly convex set is actually a spectrahedron, so we solve a potentially larger class of problems. It seems that no efficient general methods to solve hyperbolic programs are available yet. Second, even if a set is known to be a spectrahedron, it might be hard to find an explicit linear matrix inequality defining it. Third, even if the linear matrix inequality is explicitly known, it might be of much larger size than the degree of the corresponding real zero polynomial [29, 27], whereas we work directly with this polynomial.
We will now review real zero polynomials and show that there exists a regular homotopy in the space of real zero polynomials. The special case of semidefinite programming follows as a special case (see Example 9).
3.1.1 Rigidly Convex Sets
In the following, we define and examine the notion of rigidly convex sets. Moreover, we review a procedure to approximate every rigidly convex set by a smooth version thereof. Most of the content in this subsection is also part of [22, Section 2].
Definition 7 (Real zero polynomials and rigidly convex sets).
A polynomial $p\colon\mathbb{R}^{n}\to\mathbb{R}$ is called a real zero polynomial if $p(\mathbf{0})>0$ and for every $\mathbf{0}\neq\mathbf{x}\in\mathbb{R}^{n}$, the univariate polynomial
$p_{\mathbf{x}}(s)\coloneqq p(s\mathbf{x})$
has only real roots. We call the set
$\mathcal{R}(p)\coloneqq\{\mathbf{x}\in\mathbb{R}^{n}\mid p_{\mathbf{x}}\text{ has no roots in }[0,1)\}$
the rigidly convex set defined by $p$.
We call a real zero polynomial smooth, if all $p_{\mathbf{x}}$ have only roots of multiplicity $1$. This implies that $\nabla p(\mathbf{x})\neq 0$ whenever $p(\mathbf{x})=0$. Hence, the rigidly convex set $\mathcal{R}(p)$ of a smooth real zero polynomial is smooth.
In the following we present some known results about real zero polynomials, necessary for the construction of the homotopy.
Lemma 8.
Let $p\colon\mathbb{R}^{n}\to\mathbb{R}$ be a real zero polynomial. Then the following holds:
  1. (i)
    $\mathcal{R}(p)$ is closed convex with $\mathbf{0}$ in its interior.
  2. (ii)
    If $p$ is smooth and $\mathcal{R}(p)$ is bounded, then $p$ is strictly quasi-concave on $\mathcal{R}(p)$.
Proof.
For $(i)$, see [10], for $(ii)$, see [24, Lemma 2.4]. ∎
We now discuss the example of determinantal polynomials. The corresponding rigidly convex sets in this case are spectrahedra.
Example 9.
Let $A(\mathbf{x})=I_{s}+x_{1}A_{1}+\cdots+x_{n}A_{n}$ be a linear matrix polynomial, and define the polynomial
$p(\mathbf{x})\coloneqq\det(A(\mathbf{x})).$
Then $\lambda\neq 0$ is an eigenvalue of $A(\mathbf{x})$ with multiplicity $k$ if and only if $\frac{1}{1-\lambda}$ is a root of $p_{\mathbf{x}}$ with multiplicity $k$. This implies that $p$ is a real zero polynomial and $\mathcal{R}(p)=\mathcal{S}(A)$ is a spectrahedron.
3.1.2 Smoothing of Real Zero Polynomials
We now introduce the smoothing operator $\mathcal{F}_{\varepsilon}$ on the space of real-zero polynomials, that converts every rigidly convex set into a smooth rigidly convex set that approximates the original set. For this purpose, we introduce the Renegar derivative.
Definition 10 (Homogenization and the Renegar derivative).
Let $p\colon\mathbb{R}^{n}\to\mathbb{R}$ be a polynomial of degree $d$. We define the homogenization of $p$ by
$\widetilde{p}(x_{0},\mathbf{x})\coloneqq x_{0}^{d}\cdot p\left(\frac{\mathbf{x}}{x_{0}}\right).$
The Renegar derivative of $p$ is defined as
$(\partial_{R}p)(\mathbf{x})\coloneqq\left(\frac{\partial}{\partial x_{0}}\widetilde{p}\right)(1,\mathbf{x}).$
Definition 11 (Smoothing of real zero polynomials).
Let $p\colon\mathbb{R}^{n}\to\mathbb{R}$ be a real zero polynomial. We define
$\mathcal{F}_{\varepsilon}[p]\coloneqq p+\varepsilon\partial_{R}p$
as the $\varepsilon$-smoothed version of $p$.
We now state some properties of the smoothing operator.
Lemma 12.
Let $p\colon\mathbb{R}^{n}\to\mathbb{R}$ be a real zero polynomial. For $\varepsilon>0$, the $\varepsilon$-smoothed version of $p$ has the following properties:
  1. (i)
    $\mathcal{F}_{\varepsilon}[p]$ is a real zero polynomial, and hence $\mathcal{R}(\mathcal{F}_{\varepsilon}[p])$ is convex.
  2. (ii)
    If $\mathcal{R}(p)$ is bounded, then $\mathcal{R}(\mathcal{F}_{\varepsilon}[p])$ is bounded.
  3. (iii)
    If $\lambda$ is a root of multiplicity $k\geqslant 2$ of $p_{\mathbf{x}}$, then $\lambda$ is a root of multiplicity $k-1$ of $\mathcal{F}_{\varepsilon}[p]_{\mathbf{x}}$. All other roots of $\mathcal{F}_{\varepsilon}[p]_{\mathbf{x}}$ have with multiplicity $1$.
Proof.
(i) and (iii) follow from the fact that for every polynomial $h$ (with $h(\mathbf{0})\neq 0$) which has only real zeros, the polynomial $h+th^{\prime}$ also has only real zeros, with multiplicities as required. This can be deduced from general results in [4, 26] for example, or checked directly with standard calculus methods.
For (ii), let $q$ be the homogenization for $\mathcal{F}_{\varepsilon}[p]$. Then, we have that
$q(x_{0},\mathbf{x})=\widetilde{p}(x_{0},\mathbf{x})+\varepsilon x_{0}\partial_{x_{0}}\widetilde{p}(x_{0},\mathbf{x}).$
$\mathcal{R}(p)$ is bounded, if for every $\mathbf{x}\in\mathbb{R}^{n}$, there exists some $x_{0}>0$ such that $\widetilde{p}(x_{0},\mathbf{x})=0$. Similarly, $\mathcal{R}(\mathcal{F}_{\varepsilon}[p])$ is bounded, if for every $\mathbf{x}\in\mathbb{R}^{n}$, there is a $x_{0}>0$ such that $q(x_{0},\mathbf{x})=0$. Assuming the former, the latter is true, since if $h$ has a positive root, then the same is true for $h+\varepsilon th^{\prime}$, by the intermediate value theorem. ∎
3.1.3 A Regular Homotopy for Real Zero Polynomials
We now define the homotopy for optimization on rigidly convex sets. Note that the product of two real zero polynomials $p,q$ is again a real zero polynomial, and we have
$\mathcal{R}(p)\cap\mathcal{R}(q)=\mathcal{R}(pq).$ (8)
Definition 13.
Let $p_{0},p_{1}$ be smooth real zero polynomials with bounded rigidly convex sets $\mathcal{R}(p_{0})$ and $\mathcal{R}(p_{1})$. Moreover, let $\varepsilon\colon[0,1]\to[0,\infty)$ be a smooth function with $\varepsilon(0)=\varepsilon(1)=0$ and $\varepsilon(t)>0$ for $t\in(0,1)$. We define the homotopies
  1. (i)
    $\hat{p}_{t}(\mathbf{x})\coloneqq p_{0}((1-t)\mathbf{x})\cdot p_{1}(t\mathbf{x})$
  2. (ii)
    $p_{t}(\mathbf{x})\coloneqq\mathcal{F}_{\varepsilon(t)}[\hat{p}_{t}](\mathbf{x}).$
Proposition 14.
The homotopies $\hat{p}_{t}$ and $p_{t}$ satisfy the following:
  1. (i)
    If $\mathcal{R}(p_{0})$ and $\mathcal{R}(p_{1})$ are bounded, then $\mathcal{R}(\hat{p}_{t})$ and $\mathcal{R}(p_{t})$ are bounded.
  2. (ii)
    If $p_{0}$ and $p_{1}$ are smooth, then $p_{t}$ is smooth.
Proof.
(i) The boundedness of $\mathcal{R}(\hat{p}_{t})$ follows immediately from (8) and the boundedness of $\mathcal{R}(p_{t})$ from Lemma 12 (ii).
(ii) If $p_{0}$ and $p_{1}$ are smooth, we have that $s\mapsto\hat{p}_{t}(s\mathbf{x})$ has zeros of multiplicity at most $2,$ for every $t\in(0,1)$ and $0\neq\mathbf{x}\in\mathbb{R}^{n}$. Hence, applying the smoothing operator $\mathcal{F}_{\varepsilon(t)}$ gives rise to a smooth polynomial $p_{t}$ for every $t\in[0,1]$. ∎
Note that $\hat{p}_{t}$ is in general not smooth, even if $p_{0}$ and $p_{1}$ are smooth (see for example Fig. 2 for $\varepsilon=0$ and $t=0.6$)
Figure 2: Examples of the homotopy within the space of real zero polynomials for different smoothing parameters. For $\varepsilon=0$, the homotopy leads to non-smooth convex sets (see for example $t=0.5$ and $t=0.6$), as soon as $\varepsilon>0$, the roots of multiplicity $2$ become roots of multiplicity $1$.
Theorem 15.
If $p_{0},p_{1}$ are smooth and $\mathcal{R}(p_{0}),\mathcal{R}(p_{1})$ are bounded, then the homotopy $p_{t}$ from Definition 13 is regular, and therefore the induced differential equation (6) has a unique solution.
Proof.
Regularity follows from smoothness of the $p_{t}$. Indeed, if a point leaves the innermost ring of zeros of $p_{t}$ over time, this would imply a multiple zero of some $p_{t}$ on some line. ∎
Remark 16.
A similar statement to Proposition 14 can also be shown when dropping the smoothness-assumption for $\mathcal{R}(p_{1})$. In this situation $s\mapsto\hat{p}_{t}(s\mathbf{a})$ can have zeros up to multiplicity $d\coloneqq\deg(p_{1})$. Applying $\mathcal{F}_{\varepsilon(t)}$ $d$ times to $\hat{p},$ instead of only once, gives rise to a homotopy that satisfies the assumptions.
Figure 3: Optimizing a linear function on a random spectrahedron.
3.2 Convex Optimization Problems with a Single Constraint
In the following, we consider a convex optimization problem, where the non-empty feasibility set is defined by a single concave constraint. Without loss of generality, we assume that all feasible sets contain $\mathbf{0}$ in their interior.
More precisely, let $p_{0},p_{1}\colon\mathbb{R}^{n}\to\mathbb{R}$ be continuously differentiable concave functions which are strictly positive at $\mathbf{0}$. The two corresponding convex sets are defined as
$C_{i}:=\{\mathbf{x}\in\mathbb{R}^{n}\mid p_{i}(\mathbf{x})\geqslant 0\}.$
By concavity of $p_{i}$, these are closed convex sets. Moreover, assume that the objective function $f\colon\mathbb{R}^{n}\to\mathbb{R}$ is linear. We also assume to know the optimal value of
$\max_{\mathbf{x}\in C_{0}}\ f(\mathbf{x})$
and we now want to compute the optimal value for the same problem on $C_{1}$.
For this purpose, consider the homotopy
$p(t,\mathbf{x})=p_{t}(\mathbf{x})\coloneqq(1-t)p_{0}(\mathbf{x})+tp_{1}(\mathbf{x})$
for $t\in[0,1],$ which give rise to the sets
$C_{t}\coloneqq\{\mathbf{x}\in\mathbb{R}^{n}\mid p_{t}(\mathbf{x})\geqslant 0\}.$
Since $p_{t}$ is concave for every $t\in[0,1]$, every $C_{t}$ is a closed convex set. We now show that the homotopy satisfies the conditions for our method.
Proposition 17.
Let $p_{0},p_{1}$ be concave functions such that $p_{0}(\mathbf{0}),p_{1}(\mathbf{0})>0$. Then for every $t\in[0,1]$ we have:
  1. (i)
    $\mathbf{x}\in\partial C_{t}$ if and only if $p_{t}(\mathbf{x})=0$.
  2. (ii)
    $C_{t}$ is smooth for every $t\in[0,1]$, i.e. $\nabla p_{t}\neq 0$ on $\partial C_{t}$.
  3. (iii)
    $C_{0}\cap C_{1}\subseteq C_{t}\subseteq C_{0}\cup C_{1}$.
  4. (iv)
    If $p_{0}$ is strictly concave, then $p_{t}$ is strictly concave for every $t\in[0,1)$.
In particular, (iii) implies that $C_{t}$ has non-empty interior for every $t\in[0,1]$. Moreover, $C_{t}$ is bounded if $C_{0},C_{1}$ are bounded.
Proof.
(i) The only if direction is clear due to continuity. Now assume that $p_{t}(\mathbf{x})=0$. By concavity, we have that $\nabla p_{t}(\mathbf{x})\neq 0$. But this implies that for every $\varepsilon$-neighborhood of $\mathbf{x}$ there are points $\mathbf{y}$ and $\mathbf{z}$ such that $p_{t}(\mathbf{y})>0$ and $p_{t}(\mathbf{z})<0$ which shows the statement.
(ii) Using the same argument as in (i), we have that for every $t\in[0,1]$, $\nabla p_{t}(\mathbf{x})\neq 0$ for every $\mathbf{x}\in\partial C_{t}$, which shows the statement.
(iii) Let $\mathbf{x}\in C_{0}\cap C_{1}$, then $p_{0}(\mathbf{x}),p_{1}(\mathbf{x})\geqslant 0$ and therefore
$(1-t)p_{0}(\mathbf{x})+tp_{1}(\mathbf{x})\geqslant 0,$
which shows the one inclusion. For the second inclusion, assume that ${\mathbf{x}\notin C_{0}\cup C_{1}}$. Then $p_{0}(\mathbf{x}),p_{1}(\mathbf{x})<0$ and hence
$(1-t)p_{0}(\mathbf{x})+tp_{1}(\mathbf{x})<0,$
which shows the reverse inclusion.
(iv) Immediate, since the sum of a concave and a strictly concave function is strictly concave. ∎
Now let $p_{0},p_{1}$ be three times continuously differentiable concave functions, with $p_{0}$ strictly concave and $p_{0}(\mathbf{0}),p_{1}(\mathbf{0})>0$. Further assume the sets $C_{i}=\{\mathbf{x}\mid p_{i}(\mathbf{x})\geqslant 0\}$ are bounded for $i=0,1$. Then Proposition 17 together with Theorem 6 guarantees that the optimal solutions of the problems
$\max_{\mathbf{x}\in C_{t}}\ f(\mathbf{x})$
can be found by solving the differential equation (6), where the initial value is given by the unique optimal value of the problem
$\max_{\mathbf{x}\in C_{0}}\ f(\mathbf{x}).$
If $p_{1}$ is in addition strictly quasi-concave on the boundary of its nonnegativity set, then the solution of the optimization problem for $t=1$ is unique.
Example 18.
Let $p_{0}(\mathbf{x})=1-\|\mathbf{x}\|_{2}^{2}$ and $p_{1}$ a three times continuously differentiable concave function with $p_{1}(\mathbf{0})>0$ and $C_{1}$ bounded.
If $f(\mathbf{x})=\mathbf{w}^{t}\mathbf{x}$ with $0\neq\mathbf{w}\in\mathbb{R}^{n}$, the optimal point of the problem for $t=0$ is given by $\mathbf{a}_{0}=\mathbf{w}/\|\mathbf{w}\|_{2}$. Since $\nabla f=\mathbf{w}\neq 0$ on $\mathbb{R}^{n}$, we have that the solution of (6) gives rise to the optimal points of the optimization problems (2). In Fig. 4 we show a numeric example for the path of optimal solutions for
$p_{1}(\mathbf{x})=10-\sum_{k=1}^{4}\exp(\mathbf{b}_{k}^{t}\mathbf{x}),$
where $\mathbf{b}_{1}=(\frac{1}{5},\frac{1}{5})^{t}$, $\mathbf{b}_{2}=(-\frac{1}{10},\frac{1}{5})^{t}$, $\mathbf{b}_{3}=-\mathbf{b}_{1}$, and $\mathbf{b}_{4}=(\frac{1}{10},-\frac{1}{10})^{t}$, and the functional is given by $\mathbf{w}=(1,\frac{1}{2})^{t}$.
Figure 4: Homotopy of the convex optimization problem presented in Example 18.
4 Numerical Examples and Benchmarks
In the following, we present several numerical examples and compare the runtime of the homotopy method with known methods. This includes the following convex optimization problems:
  • $\triangleright$
    Optimizing a linear function on a rigidly convex set generated by the elementary symmetric polynomials (Section 4.1).
  • $\triangleright$
    Optimizing a linear function on a $k$-ellipsoid (Section 4.2).
  • $\triangleright$
    Optimizing a linear function on a $p$-norm ball (Section 4.3).
  • $\triangleright$
    Geometric programming with a single constraint (Section 4.4).
Our starting set is always the unit ball, defined by the inequality $p_{0}(\mathbf{x})=1-\|\mathbf{x}\|^{2}\geqslant 0.$ Since $p_{0}$ is both strictly concave and a real zero polynomial, it fulfills all the requirements from above. And as already mentioned in Example 18 above, computing the optimal point for a linear optimization problem on the unit ball in indeed trivial.
We emphasize that we did not focus on an optimized implementation of the homotopy method. In particular, the choice of the integrator for the differential equation is not tailored to the specific problem situation. Also, the implementation of the derivatives were done in the most basic way. Therefore, we believe that a more optimized implementation of the homotopy method might further improve the runtime of the presented method significantly.
The runtime-experiments were conducted on a standard laptop with 16GB RAM. The algorithms were run on Python, the semidefinite programs, the second-order conic programs, and the geometric programs were solved using the Mosek solver (version 9.3.20) [2] via the Cvxpy package (version 1.3.2) [1, 8]. The differential equations for the homotopy method were solved via a standard predictor-corrector method using the explicit Runge-Kutta method of order $5(4)$. The errors for the optimal points were always fixed to be at most $10^{-5}$ for all methods.
4.1 Hyperbolic Programming
In the following, we benchmark the homotopy method for hyperbolic programs (Section 3.1). Since other well-established solvers for hyperbolic programs seem not to exists, we compare our method to semidefinite programming. In particular, we consider the class of elementary symmetric polynomials $s_{k}$ (defined in (9)). While the elementary symmetric polynomial is of degree $k$, the best known way to represent its rigidly convex set needs matrices of a much larger size (see Table 1), which become intractable for SDP solvers already for small dimension $n$ and small parameters $k$.
For this purpose, let us recall the notion of hyperbolic polynomials. Let $h\colon\mathbb{R}^{n+1}\to\mathbb{R}$ be a homogeneous polynomial. We call $h$ hyperbolic in the direction $\mathbf{e}\in\mathbb{R}^{n+1}$ if the univariate polynomial $t\mapsto h(\mathbf{y}-t\mathbf{e})$ has only real roots, for every $\mathbf{y}\in\mathbb{R}^{n+1}$. Hyperbolic polynomials give rise to a closed convex cone
$\mathcal{C}_{\mathbf{e}}(h)\coloneqq\{\mathbf{y}\in\mathbb{R}^{n+1}\mid h(\mathbf{y}-t\mathbf{e})\text{ has only nonnegative roots}\},$
called the hyperbolicity cone of $h$ in direction $\mathbf{e}$ (see [10, 23, 30] for further details).
Note that hyperbolicity is precisely the homogenized version of the real zero property introduced in Section 3.1. Hyperbolic optimization is the task of optimizing a linear objective function over an affine subspace of a hyperbolicity cone. This affine subspace is precisely a rigidly convex set of a real-zero polynomial (Definition 7).
One particular class of hyperbolic polynomials are the elementary symmetric polynomials. For $n\in\mathbb{N}$ and $k\in\{1,\ldots,n\}$, let
$s_{k}(y_{0},\ldots,y_{n})\coloneqq\sum_{0\leqslant i_{1}<\ldots<i_{k}\leqslant n}y_{i_{1}}\cdots y_{i_{k}}$ (9)
be the elementary symmetric polynomial of degree $k$ in $n+1$ variables. All elementary symmetric polynomials are hyperbolic in direction $\mathbf{e}=(1,\ldots,1)$. Moreover, we have that $s_{k}(\mathbf{e})>0$ for every $k\in\{1,\ldots,n\}$, and
$\mathbb{R}_{+}^{n+1}=\mathcal{C}_{\mathbf{e}}(s_{n+1})\subseteq\mathcal{C}_{\mathbf{e}}(s_{n})\subseteq\cdots\subseteq\mathcal{C}_{\mathbf{e}}(s_{2})\subseteq\mathcal{C}_{\mathbf{e}}(s_{1})=\left\{\mathbf{y}\in\mathbb{R}^{n+1}\mid\sum_{i=0}^{n}y_{i}\geqslant 0\right\}.$
In the following, we consider a bounded affine slice of the hyperbolicity cones $\mathcal{C}_{\mathbf{e}}(s_{k})$. For this purpose, let $p_{k}$ be a dehomogenized version of $s_{k}$, defined as
$p_{k}(x_{1},\ldots,x_{n})\coloneqq s_{k}\left(e_{0}-\sum_{i=1}^{n}x_{i},e_{1}+x_{1},\ldots,e_{n}+x_{n}\right).$ (10)
Then $p_{k}$ is a real-zero polynomial, since $p_{k}(\mathbf{0})=s_{k}(\mathbf{e})>0$ and for every $s\in\mathbb{R}\setminus\{0\}$ and $\mathbf{x}\in\mathbb{R}^{n}\setminus\{0\}$, we have that
$p_{k}(\lambda\mathbf{x})=s_{k}\left(e_{0}-\sum_{i=1}^{n}\lambda x_{i},e_{1}+\lambda x_{1},\ldots,e_{n}+\lambda x_{n}\right)=\lambda^{k}s_{k}\left(\frac{1}{\lambda}\mathbf{e}+\mathbf{y}\right),$
where $\mathbf{y}=(-\sum_{i=1}^{n}x_{i},x_{1},\ldots,x_{n})$. Moreover, the rigidly convex set $\mathcal{R}(p_{k})$ is bounded for $k\geqslant 2,$ see Appendix C. Note that $\mathcal{R}(p_{k})$ corresponds (up to an isometry) to the set
$\mathcal{C}_{\mathbf{e}}(s_{k})\cap\Big{[}\mathbf{e}+\{\mathbf{x}\in\mathbb{R}^{n+1}\mid\mathbf{e}^{t}\mathbf{x}=0\}\Big{]}.$ (11)
In the following, we will benchmark the problem
$\max_{\mathbf{x}\in\mathcal{R}(p_{k})}\ f(\mathbf{x})$ (12)
where $f\colon\mathbb{R}^{n}\to\mathbb{R}$ is a linear function. We will use two different procedures:
  1. (i)
    Applying the homotopy method for hyperbolic polynomials (Section 3.1).
  2. (ii)
    Applying an SDP interior point solver to the best known spectrahedral representation of $\mathcal{R}(p_{k})$.
Specifically, we will compare the runtimes of these algorithms for the situation where $k=n-1$, as well as for the situation where $k=\lfloor\frac{n+1}{2}\rfloor$.
Regarding (ii), we use the best known spectrahedral representation of $\mathcal{C}_{\mathbf{e}}(p_{k})$ introduced in [7]. According to the construction, the matrix size of this represenation corresponds to the number of strings of length at most $k-1$ with distinct characters from an alphabet of size $n+1$. In other words, the matrix size $s(k,n)$ is given by
$s(k,n)=\sum_{i=0}^{k-1}\binom{n+1}{i}\cdot i!\geqslant\left(\frac{n+1}{k-1}\right)^{k-1}\cdot(k-1)!$
where we have used the lower bound $\binom{n}{i}\geqslant(n/i)^{i}$ for $i=k-1$. In particular, if $k\leqslant n$ and $k\sim n$, then $s(k,n)\sim c^{k}\cdot k!$ which makes the spectrahedral representation untractable already for small values $n$ and $k$. Values of $s(k,n)$ for the benchmarked situations are presented in Table 1.
$n=3$
$n=4$
$n=5$
$n=6$
$n=7$
$n=8$
$k=n-1$
$5$
$26$
$157$
$\mathbf{1100}$
$\mathbf{8801}$
$\mathbf{79210}$
$k=\lfloor\frac{n+1}{2}\rfloor$
$5$
$6$
$37$
$50$
$\mathbf{401}$
$\mathbf{5861}$
Table 1: Matrix size of the spectrahedral representation of $\mathcal{R}(p_{k})$ for different dimensions $n$. The bold cases denote the matrix sizes, where the solver did not reach a conclusion (memory error).
[…]
Figure 5: Runtimes of the homotopy method (blue) vs. the SDP solver (orange) applied to a random linear functional and $\mathcal{R}(p_{k})$ for different dimensions $n$. For the SDP solver, we use the spectrahedral representation of $\mathcal{R}(p_{k})$ from [7]. The SDP solver only reaches a conclusion for $n\leqslant 5$ and $n\leqslant 6$, respectively due to the growth of the matrix sizes (see Table 1).
Fig. 5 shows the runtimes of the homotopy method and the SDP solver for different values of $n$. While for small values of $n$, the SDP solver is much faster than the homotopy method, the SDP formulation is much slower or even intractable for larger values of $n$.
4.2 Optimizing over the $k$-Ellipsoid
In the following, we benchmark the problem of optimizing a linear function on a $k$-ellipsoid. Given points $\mathbf{u}_{1},\ldots\mathbf{u}_{k}\in\mathbb{R}^{n}$, we define the $k$-ellipsoid of dimension $n$ as the convex set
$\mathcal{E}_{n}(\mathbf{u}_{1},\ldots,\mathbf{u}_{k},r)\coloneqq\left\{\mathbf{x}\in\mathbb{R}^{n}\mid\sum_{i=1}^{k}\|\mathbf{x}-\mathbf{u}_{i}\|_{2}\leqslant r\right\}.$
The points $\mathbf{u}_{1},\ldots,\mathbf{u}_{k}$ are called the focal points of this ellipsoid.
We aim for solving the following problem:
$\begin{split}\max_{\mathbf{x}\in\mathbb{R}^{n}}&\ f(\mathbf{x})\\ \text{subject to }&\ \mathbf{x}\in\mathcal{E}_{n}(\mathbf{u}_{1},\ldots,\mathbf{u}_{k},r)\end{split}$ (13)
for different values of $n,k,$ and random choices of $\mathbf{u}_{1},\ldots,\mathbf{u}_{k}$.
In the following, we will benchmark three different methods to solve (13):
  1. (i)
    Applying our homotopy method for a single concave constraint.
  2. (ii)
    Using a spectrahedral representation of $\mathcal{E}_{n}(\mathbf{u}_{1},\ldots,\mathbf{u}_{k},r)$ to solve (13) via an SDP.
  3. (iii)
    Lifting the problem to a second-order cone program.
Regarding (i), (13) can be represented as
$\begin{split}\max_{\mathbf{x}\in\mathbb{R}^{n}}\ &f(\mathbf{x})\\ \text{subject to }&p(\mathbf{x})\geqslant 0\end{split}$ (14)
where
$p(\mathbf{x})\coloneqq r-\sum_{i=1}^{r}\|\mathbf{x}-\mathbf{u}_{i}\|_{2}.$ (15)
Note that $p$ is not differentiable at the focal points $\mathbf{u}_{i}$. Thus even if $\mathbf{u}_{i}\notin\partial\mathcal{E}_{n}(\mathbf{u}_{1},\ldots,\mathbf{u}_{k},r)$, there are times $t$ where the functions $p_{t}(\mathbf{x})$, constructed from the homotopy in Section 3.2, might not be differentiable and strictly concave. However, this happens for at most $k$ many different times. Therefore, the numerical integrator evaluates $p_{t}(\mathbf{x})$ on these problematic points with probability 00. At all other points and times, the function is concave in our above sense, so the method will work with probability $1$. The homotopy method for a $3$-ellipse is shown in Fig. 6.
Regarding (ii), in [25] it was shown that ellipsoids are spectrahedra. An explicit defining linear matrix inequality is
$\mathcal{M}(\mathbf{x})\coloneqq r\cdot I_{(n+1)^{k}}-N_{1}(\mathbf{x})\boxplus N_{2}(\mathbf{x})\boxplus\cdots\boxplus N_{k}(\mathbf{x})\succcurlyeq 0$ (16)
where
$A_{1}\boxplus A_{2}\boxplus\cdots\boxplus A_{k}\coloneqq A_{1}\otimes I\otimes\cdots\otimes I+I\otimes A_{2}\otimes I\otimes\cdots\otimes I+\ldots+I\otimes\ldots\otimes I\otimes A_{k}$
with $I$ being short-hand for the identity matrix of size $n+1$, and
$N_{j}(\mathbf{x})\coloneqq\begin{bmatrix}0&u_{j1}-x_{1}&\ldots&u_{jn}-x_{n}\\ u_{j1}-x_{1}&&&\\ \vdots&&0&\\ u_{jn}-x_{n}&&&\end{bmatrix}.$
Note that (16) is a linear matrix inequality whose matrix size is exponential in the number of focal points $k$. There is also no significantly more efficient way of defining the $k$-ellipse by a linear matrix inequality, since the degree of the defining polynomial also increases exponentially with $k$. More precisely, this degree is $2^{k}$ if $k$ is odd and $2^{k}-{k\choose k/2}$ if $k$ is even [25]. Fig. 6 shows the zero set of the real zero polynomial $\det(\mathcal{M}(\mathbf{x}))$ for a $3$-ellipse.
[…]
Figure 6: The $3$-ellipse generated by the points $\mathbf{u}_{1,2}=(\pm 2,0)$ and $\mathbf{u}_{3}=(6,0)$ with $r=12$.
Therefore (13) is equivalent to the SDP
$\begin{split}\max_{\mathbf{x}\in\mathbb{R}^{n}}\ &f(\mathbf{x})\\ \text{subject to }&\mathcal{M}(\mathbf{x})\succcurlyeq 0.\end{split}$ (17)
Regarding (iii), (13) is equivalent to second-order cone program
$\begin{split}\max_{\mathbf{x}\in\mathbb{R}^{n},\ \mathbf{s}\in\mathbb{R}^{k}}\ &f(\mathbf{x})\\ \text{subject to }\ &\sum_{i=1}^{k}s_{i}\leqslant r\\ &\|\mathbf{x}-\mathbf{u}_{i}\|_{2}\leqslant s_{i}\quad\text{ for all }i\in\{1,\ldots,k\}.\end{split}$
The results of the benchmarks are summarized in Table 2. Since the dimension of the linear matrix inequality for the $k$-ellipse increases exponentially with the number of focal points $k$, the optimization problems becomes already intractable for small $k$. In contrast, the homotopy method remains tractable also for choices of large $k$, since the description only changes by adding terms to the defining equation (15).
$k=2$
$k=3$
$k=4$
$k=5$
$k=10$
$k=100$
$n=2$ Homotopy
$0.02(1)$
$0.02(1)$
$0.06(4)$
$0.04(1)$
$0.09(4)$
$1.5(4)$
SDP
$0.03(1)$
$0.07(1)$
$3.3(4)$
$\approx 1500$*
$2^{\text{nd}}$ order cone
$0.02(1)$
$0.02(1)$
$0.03(1)$
$0.03(2)$
$0.04(2)$
$0.15(3)$
$n=4$ Homotopy
$0.05(2)$
$0.05(1)$
$0.07(3)$
$0.10(2)$
$0.21(5)$
$4(1)$
SDP
$0.06(1)$
$23(2)$
$2^{\text{nd}}$ order cone
$0.02(1)$
$0.03(1)$
$0.03(1)$
$0.03(1)$
$0.05(2)$
$0.22(3)$
Table 2: Average runtime in seconds (and the standard deviation in brackets) of optimizing a linear function for random $k$-ellipsoids in dimension $n$ with $r=2k$. The sample size is $100$.
- : Solver did not reach a conclusion (Memory error)
* : Only a single run.
4.3 Optimizing over a $p$-Norm Ball
In the following, we consider the optimization problem
$\max_{\mathbf{x}\in B_{p,r}}\ f(\mathbf{x})$ (18)
where $B_{p,r}$ is the ball of radius $r>0$ around $\mathbf{0}$ with respect to the $p$-norm, i.e.
$B_{p,r}=\left\{\mathbf{x}\in\mathbb{R}^{n}\mid s_{p}(\mathbf{x})\coloneqq r^{p}-\sum_{i=1}^{n}|x_{i}|^{p}\geqslant 0\right\}.$
For simplicity, we restrict our analysis to the case $p=8$.
Since $s_{p}$ is a convex function on $\mathbb{R}^{n}$, the homotopy introduced in Section 3.2 is a valid homotopy for this optimization problem. We will compare this method with the standard techniques via lifting (18) to a semidefinite program or to a second-order cone program.
For the SDP representation, note that $B_{8,r}$ is not a spectrahedron (see [23, Example 2.29] for a similar argument for $B_{4,r}$). However, it can be represented as the projection of a spectrahedron. This leads to the following SDP which is equivalent to (18) for $p=8$.
$\max_{\mathbf{x},\mathbf{w},\mathbf{u}\in\mathbb{R}^{n}}$ $\ f(\mathbf{x})$
subject to $\ \begin{bmatrix}r^{8}&u_{1}&\cdots&u_{n}\\ u_{1}&1&&\\ \vdots&&\ddots&\\ u_{n}&&&1\end{bmatrix}\succcurlyeq 0,$
$\ \begin{bmatrix}w_{1}&x_{1}\\ x_{1}&1\end{bmatrix}\succcurlyeq 0,\ldots,\begin{bmatrix}w_{n}&x_{n}\\ x_{n}&1\end{bmatrix}\succcurlyeq 0,$
$\ \begin{bmatrix}u_{1}&w_{1}\\ w_{1}&1\end{bmatrix}\succcurlyeq 0,\ldots,\begin{bmatrix}u_{n}&w_{n}\\ w_{n}&1\end{bmatrix}\succcurlyeq 0.$
Alternatively, $B_{8,r}$ can also be described via a projection of second-order cone constraints, i.e. constraints of the form
$\|A\mathbf{x}+\mathbf{b}\|_{2}\leqslant\mathbf{b}^{t}\mathbf{x}+d.$
More precisely, (18) for $p=8$ is equivalent to
$\max_{\mathbf{x},\mathbf{w},\mathbf{u},\mathbf{s}\in\mathbb{R}^{n}}$ $\ f(\mathbf{x})$
subject to $\sum_{i=1}^{n}s_{i}\leqslant r^{8}$
[…]
Hence, solving (18) for $p=8$ in dimension $n$ can be done in the three following ways:
  1. (i)
    Using our homotopy method of Section 3.2, i.e. solving a differential equation in $n$ dimensions.
  2. (ii)
    Solving an SDP in $\mathbb{R}^{3n}$ with matrix size $s=5n+1,$
  3. (iii)
    Solving a second-order cone program in $\mathbb{R}^{4n}$.
Note that while the homotopy method applies uniformly to all values of $p>1$, the SDP and second-order cone lifts highly depend on the choice of $p,$ which makes them more complex to implement. In particular, if $p\in\mathbb{Q}$, the dimensions of the lifts grow with the complexity of representing the fraction.
In Fig. 7 we present the run-times of these three methods for different dimensions $n$, and a random linear functions. While the homotopy method is slower than the second-order cone lift, it outperforms the SDP lift for every dimension. We expect that the gap between second-order cone programming and the homotopy method shrinks for other values of $p$, in particular those, whose second-order cone lift needs a larger overhead.
Figure 7: Solving the optimization problem (18) for $p=8$ via an SDP lift, the homotopy method, and via a second-order cone lift.
Finally note that every convex semialgebraic set of the form
$\{\mathbf{x}\in\mathbb{R}^{n}\mid p(\mathbf{x})\geqslant 0\}$
is a projected spectrahedron if $p$ is strictly quasi-concave on the set [14]. However, the matrix size $s$ and the space dimension are in general exponential in $n,$ which gives rise to even a larger gap between the runtimes of the SDP solver and the homotopy method in general.
Yet another alternative way is to compute the solutions of the Lagrange equations corresponding to the problem (18) for $p=8$, i.e. computing solutions of the system of polynomial equations:
$r^{8}$ $=\sum_{i=1}^{n}x_{i}^{8}$
$f(\mathbf{e}_{i})$ $=-8\lambda x_{i}^{7}\quad\text{ for }\ i=1,\ldots,n$
According to Bezout’s theorem, this system of equations has at most $8^{n+1}$ solutions. So the established homotopy method from numerical algebraic geometry has to track an exponential number of solution paths. Therefore, the runtime of the homotopy method for algebraic equations [11] scales exponentially in the dimension $n$. In contrast, our method tracks precisely one path, namely the path of optimal solutions.
4.4 Geometric Programming with a Single Constraint
A geometric program is a optimization problem of the following form:
$\min_{\mathbf{y}\in\mathbb{R}^{n}}$ $f_{0}(\mathbf{y})$
[…]
where $f_{0},\ldots,f_{m}$ are posynomial functions, i.e.
$f_{i}(\mathbf{y})=\sum_{k=1}^{r}c_{k}y_{1}^{b_{k1}^{(i)}}\cdots y_{n}^{b_{kn}^{(i)}}$
with $c_{k}\geqslant 0,b_{kj}^{(i)}\in\mathbb{R},$ and $r$ is the number of monomials involved in the constraint functions. By substituting $y_{j}=\exp(x_{j})$, the geometric program translates into the convex optimization problem
$\begin{split}\min_{\mathbf{x}\in\mathbb{R}^{n}}\ &g_{0}(\mathbf{x})\\ \textrm{subject to }&g_{i}(\mathbf{x})\leqslant 1\textrm{ for }i=1,\ldots,m\end{split}$ (19)
where
$g_{i}(\mathbf{x})\coloneqq\sum_{k}^{r}c_{k}\exp(\mathbf{b}_{k}^{(i)^{t}}\mathbf{x}).$
In the following, we benchmark the subclass of geometric programs with $f_{0}$ a single monomial and only one constraint, i.e. $m=1$. This is then equivalent to the problem
$\begin{split}\max_{\mathbf{x}\in\mathbb{R}^{n}}\ &f(\mathbf{x})\\ \textrm{subject to }&p(\mathbf{x})\geqslant 0\end{split}$ (20)
where $f$ is a linear function and
$p(\mathbf{x})=1-\sum_{k=1}^{r}\exp(\mathbf{b}_{k}^{t}\mathbf{x}+a_{k})$
is a concave function. Therefore, this problem can be solved by our method using the homotopy in Section 3.2.
In the benchmark, we compare our homotopy method with the interior point method for exponential cones provided by the Cvxpy package (version 1.3.2) in Python for different choices of $r$ and $n$. The plots of the runtimes are shown in Fig. 8. As the plots show, the interior point method performs better than the homotopy method for these examples. However, we believe that the homotopy method can be improved by using more tailored solvers for the differential equation, or a different homotopy which is tailored to geometric programs.
[…]
Figure 8: Runtimes of the homotopy method vs. the interior point method applied to random geometric programs. For simplicity, we choose $a_{k}=0$ and $\mathbf{b}_{k}$ to be random vectors with entries in $[-1,1]$. Moreover, we choose $C\in[r,2r]$ randomly to guarantee that 00 is in the set of feasable points. For each configuration we have a sample size of $10$. The line shows the mean runtime together with the standard deviation.
5 Conclusion and Open Questions
In this paper, we have introduced a method to solve certain classes of convex optimization problems, by introducing a homotopy between a trivial optimization problem and the target problem, and following the path of optimal solutions along the homotopy of optimization problems. In Section 2 we have shown that the path of optimal solutions is described by the non-linear differential equation (6), which has a unique solution under certain assumptions on the homotopy (Theorem 6). We have shown that this approach can be applied to examples including semidefinite and hyperbolic programs (Section 3.1) as well as convex optimization problems with a single constraint (Section 3.2). Moreover, we have shown that our method leads to a significant speed-up in certain numerical examples, most notably hyperbolic programming (Section 4.1).
The general applicability of this method begs the question under which assumptions we can also solve more general convex optimization problems with the homotopic approach. This includes for example:
  • $\triangleright$
    solving convex optimization problems with multiple constraints, or
  • $\triangleright$
    solving optimization problems with quasi-convex constraints.
Moreover, we leave an exact runtime analysis of this method for further work. This would include for example a comparision of the complexity of our method with interior-point methods applied to SDPs. Finally, since there are multiple canonical choices of homotopies for certain optimization classes (see Appendix A and Section 3.1 in the case of SDPs), this suggests to study whether there exist alternative choices of homotopies that lead to computationally more efficient results.
Acknowledgments
This research was funded in part by the Austrian Science Fund (FWF) [doi:10.55776/P33122]. For open access purposes, the authors have applied a CC BY public copyright license to any author accepted manuscript version arising from this submission. AK further acknowledges funding of the Austrian Academy of Sciences (ÖAW) through the DOC scholarship 26547. Moreover, AK wants to thank Hamza Fawzi, and Oisín Faust for fruitful discussions.
Appendix A A Homotopy within Spectrahedra not Satisfying the Assumptions
In the following, we present a canonical homotopy within the class of spectrahedra, which does not satisfy the conditions in Theorem 6. We nevertheless include this homotopy in this paper, since it might be the most intuitive choice for a homotopy to solve semidefinite programs, and it is interesting why this does not work out in general.
Let $A=(A_{1},\ldots,A_{n})$ be a tuple of symmetric $s\times s$ matrices. We define the spectrahedron generated by $A$ as
$\mathcal{S}(A)\coloneqq\{\mathbf{x}\in\mathbb{R}^{n}\mid A(\mathbf{x})=I_{s}+x_{1}A_{1}+\cdots+x_{n}A_{n}\succcurlyeq 0\}$
Note that all those spectrahedra $\mathcal{S}(A)$ contain $\mathbf{0}$ in their interior.
For $p_{A}(\mathbf{x})\coloneqq\det(A(\mathbf{x}))$ we have
$\mathcal{S}(A)=\{\mathbf{x}\in\mathbb{R}^{n}\mid s\mapsto p_{A}(s\mathbf{x})\textrm{ has no roots in }[0,1)\},$
and we have that $\mathbf{x}\in\partial\mathcal{S}(A)$ implies that $p_{A}(\mathbf{x})=0$.
Now let $\mathcal{S}(A),\mathcal{S}(B)$ be two spectrahedra defined by $n$-tuples of $s\times s$ matrices $A$ and $B$. We define the homotopy of matrix equations
$M(\mathbf{x},t)\coloneqq(1-t)A(\mathbf{x})+tB(\mathbf{x}).$
Setting $p_{t}(\mathbf{x})\coloneqq\det(M(\mathbf{x},t))$, we obtain a family of spectrahedra
$\mathcal{S}_{t}\coloneqq\{\mathbf{x}\in\mathbb{R}^{n}\mid s\mapsto p_{t}(s\mathbf{x})\textrm{ has no roots in }[0,1)\}=\mathcal{S}(M(\cdot,t)).$
However, this family of sets does not satisfy the boundedness condition necessary for Theorem 6, even if both $\mathcal{S}(A)$ and $\mathcal{S}(B)$ are bounded, as the following example shows.
Example 19.
Let
$A(\mathbf{x})=\begin{pmatrix}1-x_{1}&x_{2}\\ x_{2}&1+x_{1}\end{pmatrix}\quad\text{ and }B(\mathbf{x})=\begin{pmatrix}1-x_{2}&x_{1}\\ x_{1}&1+x_{2}\end{pmatrix}$
We have that both $p_{A}(\mathbf{x})=p_{B}(\mathbf{x})=1-x_{1}^{2}-x_{2}^{2}$, and thus
$\mathcal{S}(A)=\mathcal{S}(B)=\{\mathbf{x}\in\mathbb{R}^{2}\mid 1-x_{1}^{2}-x_{2}^{2}\geqslant 0\}$
are both the unit disk. Note that we have
$M\left(\mathbf{x},\frac{1}{2}\right)=\begin{pmatrix}1-\frac{x_{1}+x_{2}}{2}&\frac{x_{1}+x_{2}}{2}\\ \frac{x_{1}+x_{2}}{2}&1+\frac{x_{1}+x_{2}}{2}\end{pmatrix}$
and hence
$\mathcal{S}_{\frac{1}{2}}=\big{\{}\mathbf{x}\in\mathbb{R}^{2}\mid|x_{1}+x_{2}|\leqslant\sqrt{2}\big{\}},$
which is an unbounded diagonal strip. We refer to Fig. 9 for a visualization of this homotopy.
Figure 9: A visualization of the spectrahedra arising from the homotopy of matrix equations in Example 19. Since the resulting spectrahedra are not bounded, this is not a valid homotopy for our method.
Appendix B Non-Linear Objective Function and a Homotopy of Objective Functions
In the following, we generalize the homotopy method in two directions:
  • $\triangleright$
    We allow for concave objective functions instead of only linear ones.
  • $\triangleright$
    We allow for a homotopy of objective functions.
More precisely, we set up a differential equation whose solution $\mathbf{a}\colon[0,1]\to\mathbb{R}^{n}$ is the optimal value of
$\max_{\mathbf{x}\in C_{t}}\ f_{t}(\mathbf{x})$ (21)
for every $t\in[0,1],$ if the initial value $\mathbf{a}_{0}$ is the optimal value of the optimization problem for $t=0$.
We again assume that $C_{t}$ is a compact convex set for every $t\in[0,1]$, and we assume to have a regular homotopy $p$ of smooth descriptions, just as before. Moreover, we assume that $(t,\mathbf{x})\mapsto f_{t}(\mathbf{x})$ is three times continuously differentiable, that $f_{t}$ is concave, and $\nabla f_{t}\neq 0$ on $C_{t}$, i.e. all stationary points are outside of $C_{t}$. This guarantees that the optimal solution $\mathbf{a}_{t}$ satisfies $\mathbf{a}_{t}\in\partial C_{t}$.
Using Lagrange multipliers, one analogously obtains that $\mathbf{a}_{t}$ is a stationary point of the constraint problem at time $t$ if and only if
  1. ($\text{C}_{1}$)
    $\nabla f_{t}(\mathbf{a}_{t})=\lambda_{t}\nabla p_{t}(\mathbf{a}_{t})$
  2. ($\text{C}_{2}$)
    $p_{t}(\mathbf{a}_{t})=0$
where $\lambda_{t}\in\mathbb{R}$ is the Lagrange multiplier for the constraint $p_{t}$. If $f_{t}$ has its global maximum outside of $C_{t}$, we have that $\lambda_{t}<0$ (see the proof of Lemma 20).
We translate these two conditions again into an initial value problem, whose solution is a path $\mathbf{a}\colon[0,1]\to\mathbb{R}^{n};t\mapsto\mathbf{a}_{t}$ where $\mathbf{a}_{t}$ is a stationary point of the constrained optimization problem at time $t$.
Differentiating (5) and condition ($\text{C}_{2}$) using Lemma 4, we obtain again the initial value problem for stationary points $\mathbf{a}\colon[0,1]\to\mathbb{R}^{n}$:
$K(\mathbf{a}_{t},t)\cdot\frac{\mathrm{d}\mathbf{a}_{t}}{\mathrm{d}t}+m(\mathbf{a}_{t},t)=0$ (22)
where $K$ and $m$ are defined in Eq. 7. Again, the parameter $k\in\{1,\ldots,n\}$ in (7) is chosen such that $\partial_{k}f(\mathbf{a}_{t})\neq 0$ for the optimal path $\mathbf{a}_{t}$. [^7] Depending on the objective function, such a choice of $k$ might not be obvious, and one might change the choice during the time evolution.
In summary, for every solution $\mathbf{a}\colon t\mapsto\mathbf{a}_{t}$, we have that $\mathbf{a}_{t}$ is a stationary point of the constraint optimization problem of $p_{t}$ and $f_{t}$.
We now prove that there is a unique solution for the optimization problem, if $p_{t}$ and $f_{t}$ are (strictly) concave functions. For this purpose, we show that $K(\mathbf{a}_{t},t)$ is invertible. The uniqueness then follows as in Theorem 6.
Lemma 20.
Let $\mathbf{a}_{t}$ be a stationary point of the constrained problem at time $t$. Assume the following:
  1. (i)
    $p_{t}$ and $f_{t}$ are concave on $C_{t}$ where one of them is strict. [^8] The statement can also be relaxed to strict quasi-concave functions instead of strict concave functions.
  2. (ii)
    $\nabla f_{t}\neq 0$ on $C_{t}$
Then $K(\mathbf{a}_{t},t)$ is invertible.
Proof.
We start with an observation necessary for the later argument. Since $\nabla f_{t}\neq 0$ on $C_{t}$, we have that $\mathbf{a}_{t}\in\partial C_{t}$ is an optimal point of the optimization problem on $C_{t}$. Hence, by concavity of $f_{t}$, we have
$\nabla f_{t}(\mathbf{a}_{t})^{t}(\mathbf{x}-\mathbf{a}_{t})\leqslant 0$
for all $\mathbf{x}\in C_{t}$. Since $\nabla f_{t}(\mathbf{a}_{t})\neq 0$ on $C_{t}$, this implies that $\lambda_{t}<0$.
We now show that for every $\mathbf{v}\in\mathbb{R}^{n}$, we have that
$\mathbf{v}^{t}\nabla p_{t}(\mathbf{a}_{t})\neq 0\quad\text{ or }\quad\mathbf{v}^{t}\nabla\mathcal{Q}_{k,i}(\mathbf{a}_{t},t)\neq 0\quad\text{ for some }i\in\{1,\ldots,n-1\}.$
Assume that $\mathbf{v}^{t}\nabla p_{t}(\mathbf{a}_{t})=0$. This implies $\mathbf{v}^{t}\nabla f_{t}(\mathbf{a}_{t})=0$. Moreover, we have (since $\lambda_{t}<0$) and by (strict) concavity that $\mathbf{v}^{t}A(\mathbf{a}_{t},t)\mathbf{v}<0$, where
$A(\mathbf{a}_{t},t)=\nabla^{2}f_{t}(\mathbf{a}_{t})-\lambda_{t}\nabla^{2}p_{t}(\mathbf{a}_{t}).$
For every $i\in\{1,\ldots,k-1,k+1,\ldots,n\}$, we can rewrite
$\nabla\mathcal{Q}_{k,i}(\mathbf{a}_{t},t)^{t}\mathbf{v}$ $=\big{(}\nabla^{2}f_{t}(\mathbf{a}_{t})\mathbf{v}\big{)}_{k}\partial_{i}p_{t}(\mathbf{a}_{t})-\big{(}\nabla^{2}f_{t}(\mathbf{a}_{t})\mathbf{v}\big{)}_{i}\partial_{k}p_{t}(\mathbf{a}_{t})$
$+\big{(}\nabla^{2}p_{t}(\mathbf{a}_{t})\mathbf{v}\big{)}_{i}\partial_{k}f_{t}(\mathbf{a}_{t})-\big{(}\nabla^{2}p_{t}(\mathbf{a}_{t})\mathbf{v}\big{)}_{k}\partial_{i}f_{t}(\mathbf{a}_{t})$
$=\big{(}A(\mathbf{a}_{t},t)\mathbf{v}\big{)}_{k}\partial_{i}p_{t}(\mathbf{a}_{t})-\big{(}A(\mathbf{a}_{t},t)\mathbf{v}\big{)}_{i}\partial_{k}p_{t}(\mathbf{a}_{t})$
where we have used that $\partial_{i}f(\mathbf{a}_{t})=\lambda_{t}\partial_{i}p_{t}(\mathbf{a}_{t})$. By Lemma 4, $\nabla\mathcal{Q}_{k,i}(\mathbf{a}_{t},t)^{t}\mathbf{v}=0$ for all $i\in\{1,\ldots,n-1\}$ if and only if $A(\mathbf{a}_{t},t)\mathbf{v}$ is parallel to $\nabla p_{t}(\mathbf{a}_{t},t)$. Since we have
$\mathbf{v}^{t}A(\mathbf{a}_{t},t)\mathbf{v}<0\quad\text{ and }\mathbf{v}^{t}\nabla p_{t}(\mathbf{a}_{t})=0,$
this is a contradiction. ∎
Corollary 21.
Assume the same as in Theorem 6, and further that $f_{t}$ is concave, such that for some $k$ we have $\partial_{k}f_{t}\neq 0$ on $\mathbb{R}^{n}$ for every $t\in[0,1]$. Moreover, let $\mathbf{a}_{0}$ be the optimal value of
$\max_{\mathbf{x}\in C_{0}}\ f_{0}(\mathbf{x})$
Then the unique solution of (22) with initial value $\mathbf{a}_{0}$ for $t=0$ corresponds to the optimal values of the optimization problems (21).
Proof.
Identical to the proof of Theorem 6. ∎
Example 22.
Let
$C_{1}=\{\mathbf{x}\in\mathbb{R}^{n}\mid p_{1}(\mathbf{x})\geqslant 0\},$
where $p_{1}$ is a three times differentiable, concave function. In the following, we want to compute the closest point/distance of a point $\mathbf{y}\notin C_{1}$ to $C_{1}$. This corresponds to the following optimization problem:
$\max_{\mathbf{x}\in\mathbb{R}^{n}}$ $-\|\mathbf{x}-\mathbf{y}\|_{2}$
subject to $p_{1}(\mathbf{x})\geqslant 0.$
Let $p_{0}(\mathbf{x})=1-\|\mathbf{x}\|_{2}^{2}$. Moreover, let $f_{0}(\mathbf{x})=\mathbf{w}^{t}\mathbf{x}$ with $0\neq\mathbf{w}\in\mathbb{R}^{n}$ and ${f_{1}(\mathbf{x})=-\|\mathbf{x}-\mathbf{y}\|_{2}}$.
We consider the homotopy of objective functions
$f_{t}(\mathbf{x})=(1-t)f_{0}(\mathbf{x})+tf_{1}(\mathbf{x})$
Note that the optimal point of the problem for $t=0$ is given by $\mathbf{a}_{0}=\mathbf{w}/\|\mathbf{w}\|_{2}$. In Fig. 10 we show a numeric example for the path of optimal solutions for the concave function
$p_{1}(\mathbf{x})=5-\log\left(\sum_{k=1}^{4}\exp(\mathbf{b}_{k}^{t}\mathbf{x})\right),$
where $\mathbf{b}_{1}=(2,2)^{t}$, $\mathbf{b}_{2}=(-\frac{8}{5},\frac{8}{5})^{t}$, $\mathbf{b}_{3}=(-\frac{3}{5},\frac{6}{5})$, and $\mathbf{b}_{4}=(0,-\frac{6}{5})^{t}$. The inital linear function is given by $\mathbf{w}=(-1,1)^{t}$ and the distance is minimized to $\mathbf{y}=(-6,\frac{5}{2})^{t}\in C_{1}$.
(a) The initial optimization problem, i.e. optimizing $\mathbf{x}\mapsto x_{2}-x_{1}$ on a Euclidean ball.
(b) The target optimization problem. Computing the minimal distance of the point $(-6,2.5)$ to a convex set.
Figure 10: The homotopy method applied to the optimization problem described in Example 22. The path of optimal points is shown in both figures in grey.
Appendix C Boundedness of the Rigidly Convex Sets of Symmetric Polynomials
In the following, we prove that $\mathcal{R}(p_{k})$ (see (10)) is bounded for $k\geqslant 2$.
Proposition 23.
$\mathcal{R}(p_{k})$ is bounded for $k\geqslant 2$.
Proof.
Since $\mathcal{R}(p_{k})$ is an affine slice of the hyperbolicity cones $\mathcal{C}_{\mathbf{e}}(s_{k})$ (see (11)), we have that
$\mathcal{R}(p_{k+1})\subseteq\mathcal{R}(p_{k}),$
by [23, Remark 2.75].
Therefore it suffices to show that $\mathcal{R}(p_{2})$ is bounded. We show that for every $\mathbf{0}\neq\mathbf{a}\in\mathbb{R}^{n}$, the polynomial $\lambda\mapsto p_{2}(\lambda\mathbf{a})$ has roots with different sign. Elementary manipulations lead to
$p_{2}(\mathbf{x})=\frac{n(n+1)}{2}-\sum_{1\leqslant i\leqslant j\leqslant n}^{n}x_{i}x_{j}.$
Hence, for every $\mathbf{a}\in\mathbb{R}^{n}$, we have that
$p_{2}(\lambda\cdot\mathbf{a})=\frac{n(n+1)}{2}-b\lambda^{2}$
with
$b\coloneqq\sum_{1\leqslant i\leqslant j\leqslant n}a_{i}a_{j}>0.$
Thus the two roots must have different sign, which implies that $\mathcal{R}(p_{2})$ is bounded. ∎